OpenStudy (anonymous):

A truck enters a highway driving 60 mph. A car enters the highway at the same place 9 minutes later and drives 69 mph in the same direction. From the time the car enters the highway, how long will it take the car to pass the truck?

OpenStudy (oaktree):

If the solution is at n hours, then 60n<69n-9/60.

OpenStudy (anonymous):

solution in minutes

OpenStudy (oaktree):

9n>9/60, and n>1/60, so after one minute.

OpenStudy (anonymous):

Yes it does thanks!

OpenStudy (lgbasallote):

hmm wait...

OpenStudy (lgbasallote):

i forgot to convert...

OpenStudy (anonymous):

yeah I just got a negative number

OpenStudy (lgbasallote):

60mph = 1 mile per minute 69 mph = 1.15 mile per minute so the distance the car would have traveled when it overtakes will be d = 1.15t the distance the truck would have traveled will be d = 1(t+9) so equate them 1.15t = t + 9 solve for t

OpenStudy (anonymous):

t=60 that would mean 60 minutes right?

OpenStudy (lgbasallote):

yeah

OpenStudy (anonymous):

cool thanks!

OpenStudy (lgbasallote):

welcome